HUBBLE SPACE TELESCOPE DAILY REPORT #2444 PERIOD COVERED: 0000Z (UTC) 08/25/99 - 0000Z (UTC) 08/26/99 Daily Status Report as of 238/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD 8408 (STIS CCD Dark Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the darks for the STIS CCD. The observations were executed as scheduled, and no anomalies were reported. 1.2 Completed Three Sets of WF/PC-2 8443 (WF/PC-2 Cycle-8 Supplemental Darks pt1/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no problems were reported. 1.3 Completed WF/PC-2 8398 (Saturn's Rings) The WF/PC-2 was used to make observations of the rings around the planet Saturn. Saturn's rings are a unique dynamical laboratory and continue to present challenging problems that require additional data to attack. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed STIS/CCD 8409 (STIS CCD Bias Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and to track the evolution of hot columns. The observations were executed as scheduled, and no anomalies were reported. 1.5 Completed Three Sets of STIS/MA1 8426 (Cycle 8 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1) was used to make a routine Cycle-8 MAMA dark noise calibration measurement. This proposal will provide the primary means of checking on the health of the MAMA detectors. This is done through frequent monitoring of the background count rate. The observations were completed as planned, and no problems were reported. 1.6 Completed Two Sets of WF/PC-2/STIS/CCD 7496 (Star Formation Histories of Dwarf Irregular Galaxies II: the Old Populations of IC 1613 and Sextans A) The WF/PC-2 and the Space Telescope Imaging Spectrograph (CCD) were used to make observations of the dwarf irregular galaxy IC1613. Recent results from the ground and from HST for different types of dwarf galaxies show that even in these apparently simple stellar systems, a variety of different star formation histories can be found. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed WF/PC-2 8457 (UV Earthflats) The WF/PC-2 was used to make a ultraviolet Earth flat calibration image. This is to improve the quality of pipeline the ultraviolet flatfields. The observations were completed as scheduled, and no problems were reported. 1.8 Completed STIS/MA1 7762 (The Gaseous Environment of the Ursa Minor Dwarf Spheroidal Galaxy) The Space Telescope Imaging Spectrograph (MA1) was used to probe the interstellar gas content of the Ursa Minor dwarf spheroidal galaxy QJ0640-503. A wavelength calibration image was also taken. The observations completed as planned, with no reported problems. 1.10 Completed WF/PC-2 8453 (Cycle 8 Polarization) The WF/PC-2 was used to make observations of the standard source star BD+64D106 in order to verify the stability of the instrument’s polarization calibration. The observations were completed as planned, and no problems were reported. 1.11 Completed WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed Two Sets of STIS/CCD 8271 (A Snapshot Survey of LMC Planetary Nebulae: A Study of Nebular and Late Stellar Evolution) The Space Telescope Imaging Spectrograph (CCD) was used to make medium-resolution slitless spectroscopy snapshot images of the planetary nebulae LMC-SMP-71 and LMC-SMP-25 located in the Large Magellanic Cloud to study the parallel evolution of the nebulae and their central stars. The observations were completed as planned, and no problems were reported. 1.13 Completed WF/PC-2 8436 (POMS Test Proposal: WFII Backup Parallel Archive Proposal II) The WF/PC-2 was used to make archive observations parallel to the prime science observations as part of a POMS test proposal. This is a POMS test proposal designed to simulate scientific plans. The observations were completed as planned, and no anomalies were reported. 1.14 Completed WF/PC-2 8192 (A Snapshot Survey of Probable Nearby Galaxies) The WF/PC-2 was used to make a snapshot survey images of a probable nearby galaxy known as Cas-DSPH. The census of galaxies in the Local Volume of space is very incomplete, particularly at the faint end. The observations were executed as scheduled, and no anomalies were noted 1.15 Completed STIS/CCD 8406 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation V) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude observations taken in parallel with the prime science observations as part of a POMS proposal. This is a STIS non-scripted parallel Condition V proposal. The observations were completed as scheduled, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 9 Successful: 8 Scheduled Re-acquisitions: 8 Successful: 5 As documented in HSTAR 7235, the acquisition at 238/032053Z failed when the search radius limit was exceeded on FGS-3. The following re-acquisitions at 238/055819Z, 238/063510Z, and 238/081158Z also failed. Three proposals occurring during the period of tomorrow’s report were affected: STIS/MA1/MA2 7764, WF/PC-2 8436, and WF/PC-2 8444. 2.2 FHST Updates: Scheduled: 23 Successful: 23 2.3 Operations Notes: The flight operations team remains in the Building 23 back-up facility. At 237/2120Z they began using the CCS “G” string. While on that string, the generation of engineering files has not been successful, despite the intervention of the Response Team and System Administrators. A re-cycle of the SAC server did not clear the problem. HSTAR 7234 was written. Using ROP SR-1, the SSR EDAC error counter was cleared four times. Per ROP NS-3, the NSSC-1 status buffer was dumped at 237/1414Z. At 237/174954Z, near the end of orbit night, the battery 4 voltage flagged low at 27.63 volts (the limit is 27.64 volts). There was no vehicle maneuver in progress at the time. HSTAR 7233 was generated. Following the acquisition/re-acquisition problems described in 2.1, the resultant STIS flight software error counter was reset at 238/0741Z per ROP NS-12. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH